# Semantic Component Analysis: Discovering Patterns in Short Texts Beyond Topics

### v1

@@ Abstract
Topic modeling is a key method in text analysis, but existing approaches are limited by assuming one topic per document or fail to scale efficiently for large, noisy datasets of short texts. We introduce Semantic Component Analysis (SCA), a novel topic modeling technique that overcomes these limitations by discovering multiple, nuanced semantic components beyond a single topic in short texts which we accomplish by introducing a decomposition step to the clustering-based topic modeling framework. Evaluated on multiple Twitter datasets, SCA matches the state-of-the-art method BERTopic in coherence and diversity, while uncovering at least double the semantic components and maintaining a noise rate close to zero while staying scalable and effective across languages, including an underrepresented one.
@@ 1 Introduction
Topic modeling is an essential tool for automated text analysis in digital humanities, computational social science, and related disciplines Grimmer et al. (2022); Rauchfleisch et al. (2023); Blei (2012); Maurer and Nuernbergk (2024) where it aims to find high-level semantic patterns, the topics, in a corpus and then assigning them to documents. The state-of-the-art for large-scale datasets, BERTopic Grootendorst (2022), assumes a single topic per document, which often does not hold even for short texts Paul and Girju (2010), as can be seen in Figure 1. BERTopic and related approaches do thus not model a large portion of the information in the documents. Besides, while they are able to achieve high topic coherence and diversity, a high rate of samples is usually assigned to no topic, i.e. they have a high noise rate Egger and Yu (2022).
Figure 1: A Tweet from the Trump dataset decomposed into its semantic components as defined by the Top-3 component activation scores of iteration 1 (components 6, 24, and 0) and subsequent iterations (components 56, 118, 70). Component 0 refers to the most dominant topic in the Trump dataset, of which a majority of Tweets is about Donald Trump himself.
To overcome these limitations, we propose Semantic Component Analysis (SCA) as a novel method to discover multiple patterns per document, which we coin “semantic components". SCA is a topic modeling technique, that - in addition to topics - formalizes the discovery of nuanced semantic components and scales to large datasets while maintaining a noise rate close to zero and matching BERTopic in topic coherence and diversity. Based on a sequence embeddings-based clustering pipeline as proposed by Angelov (2020), for each cluster, we introduce a decomposition step and iterate over the residual embeddings. This decomposition enables us to represent a sample’s content independently of the semantic components found in previous iterations and to find additional components.
Figure 2: Illustration of the basic clustering algorithm for linear feature discovery. Top-Right: Top-4 components (by cluster size) of the first iteration, that is equivalent to a BERTopic run. Bottom-right: components uncovered by SCA in subsequent iterations (see Table 7 in the appendix for more examples).
We evaluate our method on three Twitter datasets, one of all the Tweets of Donald Trump (Brown, 2021), another one compiled of Chinese-language news outlets that we scraped ourselves and publish alongside this paper, and finally a corpus of Tweets in Hausa (Muhammad et al., 2022). We show that our method is able to match the performance of BERTopic (Grootendorst, 2022) in terms of topic coherence and diversity, while additionally providing a large number of semantic components missed by the state-of-the-art approach. We also show that our method can provide an interpretable transformation that matches the performance of the base embedding in a variety of downstream tasks.
@@
  • We propose a novel topic modeling method, Semantic Component Analysis (SCA), that can discover multiple semantic components per document.
  • We show that our method is able to match state-of-the-art topic modeling, while providing additional, more nuanced insights into the semantic patterns in the data and maintaining a noise rate close to zero.
@@ 2 Related Work
Topic modeling’s goal is to find semantic patterns in a corpus and to assign them to the documents. Latent Dirichlet Allocation (LDA) (Blei et al., 2003; Pritchard et al., 2000) works well on small datasets of long texts and has been generalized to larger datasets Hoffman et al. (2012). Modeling a distribution over topics for each document, it can account for multi-topic documents. However, it performs poorly on short texts Ushio et al. (2021); de Groot et al. (2022), is generally vulnerable to noisy data, and produces topics that lack domain adaption Egger and Yu (2022). Recent advancements have integrated neural networks (Srivastava and Sutton, 2017; Liu et al., 2015; Dieng et al., 2019; Qiang et al., 2016; Zhao et al., 2021). Topic modeling based on LLMs Pham et al. (2024); Doi et al. (2024) has also gained attention, though their cost is a barrier for most large-scale applications. Methods based on clustering embeddings stand out because of their efficiency, domain-adaption, and topic coherence on large corpora of short texts. Top2Vec (Angelov, 2020) clusters document embeddings generated by Doc2Vec (Le and Mikolov, 2014). There, UMAP (McInnes et al., 2020) reduces embedding dimensionality mitigating the “curse of dimensionality" (Reimers and Gurevych, 2021), and clustering is performed using HDBSCAN (Campello et al., 2015). BERTopic (Grootendorst, 2022) improves the Top2Vec framework by using SBERT embeddings (Reimers and Gurevych, 2019) and c-TF-IDF weighting for representing the topics. It is considered state-of-the-art for large datasets of short texts Egger and Yu (2022). However, BERTopic does not provide a way to generalize to multi-topic documents. The cluster-hierarchy of HDBSCAN could be used to assign multiple topics per sample, but this approach is insufficient for finding patterns that are not subtopics of each other. Besides, BERTopic’s results usually suffer from a high noise rate Egger and Yu (2022), which can range at up to 74% de Groot et al. (2022). In summary, the topic modeling field does not provide a solution for scenarios with large, noisy datasets of short texts containing multiple patterns of interest per sample, that matches the state-of-the-art topic quality.
@@ 3 Methodology
SCA is a clustering pipeline that introduces a novel decomposition step to find semantic components in the embedding space, which are represented by unit vectors. For clustering, we employ the cosine distance in UMAP to map the embeddings to a 5-dimensional, Euclidean space on which HDBSCAN operates, mirroring the setup in BERTopic. Visualized in Figure 2, it works as follows:
@@
    $\mathbf{x}_{i}^{\prime}=\mathbf{x}_{i}-\mu\mathbb{1}_{\alpha_{i,j}>\alpha}\langle\mathbf{x}_{j},\mathbf{v}_{i}\rangle\mathbf{v}_{i}$ (2)
@@
After each decomposition step in 3, we set $\mathbf{x}_{j}=\mathbf{x}_{j}^{\prime}$ for all $j$ and repeat it for the next cluster. The algorithm converges when the residuals are zero, but we introduce other practical stopping criteria listed in Appendix 8.3. Note, that we introduce two hyperparameters: $\mu\in[0,1]$ controlling the amount of the decomposition and a threshold $\alpha$, which is compared against $\alpha_{i,j}$, the cosine similarity between $\mathbf{x}_{j}$ and $\mathbf{v}_{i}$. See Appendix 8.1 for a detailed discussion of $\alpha$ and $\mu$. We assume that each cluster found represents a semantic component. To obtain an interpretable representation, we follow BERTopic and use class-based (c-) TF-IDF to rank tokens according to their importance for a cluster, from which we obtain the TOP-10 set of tokens for each component.
@@
We introduce a merging step that computes the token overlap between two components as the fraction of tokens their representations share. Here, we introduce a threshold $\theta\in[0,1]$ and merge two components into a multi-dimensional component if the token overlap is larger than $\theta$ and keep the representation of the component with the lower ID (i.e. the component that appeared earlier in the procedure). That way, we avoid duplicates arising in larger clusters spread across a non-linear or multi-dimensional part of the embedding space.
@@
Finally, we include the Medoid of each cluster, which is the sample closest to the centroid in terms of cosine similarity. We note, that SCA employs the same modular approach as BERTopic, which allows the embedding, dimensionality reduction, clustering, and representation method to be replaced by similar methods.111see https://maartengr.github.io/BERTopic/ Furthermore, our specific choices for these parts of the model, namely UMAP and HDBSCAN, make the first iteration of SCA equivalent to a BERTopic run, which allows for a direct comparison of the methods.
3.1 Evaluation
@@
We evaluate our method on three datasets of different sizes and languages. [^2] https://github.com/mainlp/semantic_components For an English-language use-case, we include all Tweets by Donald Trump until December 2024 (60K samples, Brown (2021)). To evaluate on a high-resource non-English language, we collected and publish 1.5M Chinese-language Tweets from a list of news outlets (details in Appendix 9.1). To cover a low-resource language, we include a set of 512K Tweets in Hausa collected by Muhammad et al. (2022). We evaluate our method in terms of number of components, noise rate, topic coherence, and diversity (see Appendix 9.2 for definitions) and compare it to the highly popular BERTopic. Furthermore, we perform a grid run over the hyperparameters to assess their influence. To check if the components found are indeed beyond what is found by BERTopic, we compute the token overlap between the components found by SCA and the topics found by BERTopic, as well as the sample overlap with the cluster hierarchy in HDBSCAN of BERTopic and the components found by SCA (see Appendix 9.2 for details). Finally, we evaluate the transformation of the embeddings into the component space in terms of its performance in downstream tasks in the MTEB benchmark (Muennighoff et al., 2023).
@@ 4 Results
We provide a table of the Top-10 components of the first three iterations of runs on each dataset in Appendix 9. A subset can be found in Figure 2. The metrics table can be found in Figure 3 while that of the grid runs can be found in Appendix 9.2. For all datasets, the method finds a large number of components, with the Trump dataset yielding 182 components in total after merging, the Chinese News dataset 594 and the Hausa dataset 331, surpassing the 55, 212, and 102 respective topics found by BERTopic. These components relate to patterns distinct from the topics found by BERTopic, which is evident from the low average maximum token overlap scores of 1.85, 1.39 and 1.32 on the three datasets respectively. On the Trump dataset, only 8 components with a token overlap of higher than 0.5 are merged, on the Hausa dataset it is 9, and for the Chinese News dataset 19. The average maximum sample overlaps with any of BERTopic’s hierarchical subtopics are small and between 0.077 and 0.093 on all datasets. This indicates that the additional components are not sub- or supertopics of the topics found by BERTopic. The noise rate is close to zero for all datasets, i.e. almost all samples are assigned at least one component. The topic coherence scores are high for all datasets, with the Chinese dataset showing the highest NPMI coherence of 0.120 and the Trump dataset showing the lowest of -0.168, all matching the baseline. The topic diversity is high for all datasets, between 0.7 and 0.8, but about 10-15% lower than in the baseline. In sum, this shows that the quality of the components found matches that of BERTopic while providing a more nuanced view of the data through the large number of extra components found and the low noise rate. Comparing the results over the hyperparameters on the Trump dataset, we find that for $\alpha>0.7$ and $\mu<0.4$, the method reliably finds a large number of semantic components at least double the number of topics found by BERTopic, usually ranging from three to four times.
@@
The additional components found relate to patterns that intuitively match the expected distribution of the datasets. For example, in the Trump dataset, beyond the topics about Trump himself, entrepreneurship, and election campaigning found by BERTopic, we uncover components about “fake news", inspiration and role models, paying condolences, and the “witch hunt" narrative among others (see Appendix 9.3) underlining SCA’s ability to find important patterns not captured by the baseline. This view is reinforced by the component distributions of single samples, as shown in Figure 1. Similarly, in the Chinese dataset, the additional components relate to important motives in Chinese news reporting, such as protests, family, the trade war with the US, and currency (see Appendix 9.4). In the Hausa dataset, the additional insights include components about poverty, courts and law, and work (see Appendix 9.5). However, we note, that in this low-resource scenario, some components relate to semantically more narrow concepts, which is possibly due to a lack of expressive power in the base embeddings.
@@
Finally, the results of the topic distribution provided by SCA can match the performance of PCA and the base embedding model in the MTEB benchmark (see Appendix 9.6), indicating the method’s faithfulness in representing the information of the samples in the component space.
@@ 5 Conclusion
We introduce Semantic Component Analysis as a novel topic modeling method that is able to discover multiple semantic components per document. We show, that the components found by SCA indeed provide new, more nuanced insights into the data while maintaining a noise rate close to zero and matching the state-of-the-art in terms of topic quality. Furthermore, we show that the method is able to generalize to multiple languages, including a low-resource African language, and provide a faithful transformation of the embeddings into the component space. We believe that SCA provides a significant improvement in large-scale text analysis and opens up new possibilities for the interpretation of semantic patterns in short texts.
6 Limitations
@@
Acknowledgments
We’re grateful to Kristin Shi-Kupfer and David Adelani for consulting on the Chinese and Hausa datasets respectively. Furthermore, we would like to thank Xinpeng Wang, Ryan Soh-Eun Shim, Beiduo Chen, and Yang Janet Liu for their valuable feedback on earlier drafts of this paper.
Appendix
7 Risks and Ethical Considerations
@@
Firstly, there is a risk of misinterpretation or misrepresentation of the discovered semantic components. The interpretation of these components relies on the assumption that they represent meaningful and coherent topics or themes in the data. However, there is always a possibility of misinterpreting or misrepresenting the true meaning of these components, leading to incorrect conclusions or biased interpretations. Even earlier in the pipeline, there is a risk of algorithmic bias in the clustering and representation process. Especially the base embedding model can introduce biases based on the training data and the model architecture that SCA can inherit. It is important to carefully select and validate thesethe base model to ensure as fair and unbiased representations as possible.
@@
Thirdly, the nature of online social media data, such as Tweets, can introduce additional risks and challenges. Social media data can be subject to manipulation or misinformation, and contain offensive or harmful content. SCA may inadvertently amplify or propagate such content if not properly filtered or evaluated. It is important to consider the potential impact of the data and the extracted semantic components on individuals, communities, and society as a whole. Furthermore, the data we collect and
Lastly, there may be ethical considerations regarding the use of SCA in certain domains or applications. For example, if SCA is used in the context of social media analysis or opinion mining, there is a risk of amplifying or reinforcing existing biases or stereotypes present in the data. It is important to critically evaluate and address these ethical considerations when applying SCA in such contexts.
@@
8 Methodology Appendix
8.1 Hyperparameter Discussion

### v2

@@ Abstract
Topic modeling is a key method in text analysis, but existing approaches are limited by assuming one topic per document or fail to scale efficiently for large, noisy datasets of short texts. We introduce Semantic Component Analysis (SCA), a novel topic modeling technique that overcomes these limitations by discovering multiple, nuanced semantic components beyond a single topic in short texts which we accomplish by introducing a decomposition step to the clustering-based topic modeling framework. We evaluate SCA on Twitter datasets in English, Hausa and Chinese. It achieves competetive coherence and diversity compared to BERTopic, while uncovering at least double the semantic components and maintaining a noise rate close to zero. Furthermore, SCA is scalable and effective across languages, including an underrepresented one.
Semantic Component Analysis:
Discovering Patterns in Short Texts Beyond Topics
Florian Eichin1, Carolin M. Schuster2, Georg Groh2, and Michael A. Hedderich1 1LMU Munich and Munich Center for Machine Learning (MCML) 2Technical University of Munich {feichin,hedderich}@cis.lmu.de, {carolin.schuster,groh}@in.tum.de
@@ 1 Introduction
Topic modeling is an essential tool for automated text analysis in digital humanities, computational social science, and related disciplines Grimmer et al. (2022); Rauchfleisch et al. (2023); Blei (2012); Maurer and Nuernbergk (2024). It aims to find high-level semantic patterns, the topics, in a corpus and then assign them to documents. The state-of-the-art for large-scale datasets, BERTopic Grootendorst (2022), assumes a single topic per document, which often does not hold even for short texts Paul and Girju (2010), as can be seen in Figure 1. BERTopic and related approaches thus do not model a large portion of the information in the documents. Additionally, while they achieve high topic coherence and diversity (defined in Appendix B.2), a high rate of samples is usually assigned to no topic, i.e. they have a high noise rate Egger and Yu (2022).
Figure 1: A Tweet from the Trump dataset (Brown, 2021) decomposed into its semantic components as defined by the Top-3 component activation scores of iteration 1 (components 6, 24, and 0) and subsequent iterations (components 56, 118, 70). Component 0 refers to the most dominant topic in the Trump dataset, of which a majority of Tweets is about Donald Trump himself.
To overcome these limitations, we propose Semantic Component Analysis (SCA) as a novel method to discover multiple patterns per document, which we coin semantic components. SCA is a topic modeling technique, that — in addition to topics — formalizes the discovery of nuanced semantic components. It scales to large datasets while maintaining a noise rate close to zero and matches BERTopic in topic coherence and diversity. Based on clustering sequence embeddings as proposed by Angelov (2020), for each cluster, we introduce a decomposition step and iterate over the residual embeddings (see Figure 2). This decomposition enables us to represent a sample’s content independently of the semantic components found in previous iterations and to find additional components.
Figure 2: Illustration of our proposed method, that clusters residual embeddings obtained from linear decompositions of the embedding space. Top-Right: Top-4 components (by cluster size) of the first iteration equivalent to a BERTopic run. Bottom-right: components uncovered in subsequent iterations (see Appendix C.2 for more).
We evaluate our method on three Twitter datasets, one of Tweets by Donald Trump (Brown, 2021), another one of Tweets in Hausa (Muhammad et al., 2022), and finally one compiled from Chinese-language news outlet accounts that we scraped ourselves and publish alongside this paper. We show that our method is able to match the performance of BERTopic in terms of topic coherence and diversity, while additionally providing a large number of semantic components missed by the state-of-the-art approach. We also show that our method can provide an interpretable transformation that matches the performance of the base embedding in a variety of downstream tasks.
@@
  • We propose a novel topic modeling method for short texts, Semantic Component Analysis (SCA), that can discover multiple patterns per sample, which we coin "semantic components".
  • We show that our method is able to match state-of-the-art topic modeling, while providing additional, more nuanced insights into the data and achieving a noise rate close to zero.
@@ 2 Related Work
Topic modeling’s goal is to find semantic patterns in a corpus and to assign them to the documents. Latent Dirichlet Allocation (LDA, Blei et al. 2003; Pritchard et al. 2000) works well on small datasets of long texts and has been generalized to larger datasets Hoffman et al. (2012). Modeling a distribution over topics for each document, it can account for multi-topic documents. However, it performs poorly on short texts Ushio et al. (2021); de Groot et al. (2022), is generally vulnerable to noisy data, and produces topics that lack domain adaption Egger and Yu (2022). Recent advancements have integrated neural networks (Srivastava and Sutton, 2017; Liu et al., 2015; Dieng et al., 2019; Qiang et al., 2016; Zhao et al., 2021). Topic modeling based on LLMs Pham et al. (2024); Doi et al. (2024) has also gained attention, though their cost is a barrier for most large-scale applications. Methods based on clustering embeddings stand out because of their efficiency, domain-adaption, and topic coherence on large corpora of short texts. Top2Vec (Angelov, 2020) clusters document embeddings generated by Doc2Vec (Le and Mikolov, 2014). There, UMAP (McInnes et al., 2020) reduces embedding dimensionality mitigating the "curse of dimensionality" (Reimers and Gurevych, 2021), and clustering is performed using HDBSCAN (Campello et al., 2015). BERTopic (Grootendorst, 2022) improves the Top2Vec framework by using SBERT embeddings (Reimers and Gurevych, 2019) and c-TF-IDF weighting for representing the topics. It is considered state-of-the-art for large datasets of short texts Egger and Yu (2022). However, BERTopic does not provide a way to generalize to multi-topic documents. The cluster-hierarchy of HDBSCAN could be used to assign multiple topics per sample, but this approach is insufficient for finding patterns that are not subtopics of each other. Additionally, BERTopic’s results usually suffer from a high noise rate Egger and Yu (2022), which can range at up to 74% de Groot et al. (2022). In summary, the topic modeling field does not provide a solution for scenarios with large, noisy datasets of short texts containing multiple patterns of interest per sample, that matches the state-of-the-art topic quality.
@@ 3 Methodology
SCA is a clustering pipeline that introduces a decomposition step to find semantic components in the embedding space, which are represented by unit vectors. For clustering, we employ the cosine distance in UMAP to map the embeddings to a 5-dimensional, Euclidean space on which HDBSCAN operates, mirroring the setup in BERTopic. Visualized in Figure 2, it works as follows:
@@
    $\mathbf{x}_{i}^{\prime}=\mathbf{x}_{i}-\mu\mathbbm{1}_{\alpha_{i,j}>\alpha}\langle\mathbf{x}_{j},\mathbf{v}_{i}\rangle\mathbf{v}_{i}$ (2)
@@
After each decomposition step in 3, we set $\mathbf{x}_{j}=\mathbf{x}_{j}^{\prime}$ for all $j$ and repeat it for the next cluster. The algorithm converges when the residuals are zero. We introduce other practical stopping criteria listed in Appendix A.3. Note, that SCA has two hyperparameters: $\mu\in[0,1]$, controlling the amount of the decomposition, and a threshold $\alpha$, which is compared against $\alpha_{i,j}$, the cosine similarity between $\mathbf{x}_{j}$ and $\mathbf{v}_{i}$. See Appendix A.1 for a detailed discussion of $\alpha$ and $\mu$. We assume that each cluster found represents a semantic component. To obtain an interpretable representation, we follow BERTopic and use c-TF-IDF to rank the tokens according to their importance for a cluster storing the Top-10 set of tokens for each cluster.
@@
We introduce a merging step that computes the token overlap between two components as the fraction of tokens their representations share. Here, we introduce a threshold $\theta\in[0,1]$ and merge two components into a multi-dimensional component if the token overlap is larger than $\theta$. We keep the representation of the component that appeared earlier in the procedure. That way, we avoid duplicates arising in larger clusters spread across multiple dimensions of the embedding space.
Finally, we include the medoid of each cluster, which is the sample closest to the centroid in terms of cosine similarity, as a representation.
@@
We note, that SCA employs the same modular approach as BERTopic, which allows the embedding, dimensionality reduction, clustering, and representation method to be replaced by similar methods.111see https://maartengr.github.io/BERTopic/ Furthermore, our specific choices for these parts of the model, namely UMAP and HDBSCAN, make the first iteration of SCA equivalent to a BERTopic run, which allows for a direct comparison of the methods.
4 Data & Evaluation
@@
We evaluate our method on three datasets of different sizes and languages.222code and data provided in the supplementary materials For an English-language use-case, we include all Tweets by Donald Trump until December 2024 (60K samples, Brown 2021). To cover a low-resource language, we include a set of 512K Tweets in Hausa collected by Muhammad et al. (2022). To evaluate on a high-resource non-English language, we collect and publish 1.5M Chinese-language Tweets from a list of news outlets (details in Appendix B.1). We evaluate our method in terms of number of components, noise rate, topic coherence, and diversity (see Appendix B.2 for definitions) and compare it to the highly popular BERTopic. Furthermore, we perform a grid run over the hyperparameters to assess their influence. To check if the components found are indeed beyond what is found by BERTopic, we compute the token overlap between the components found by SCA and the topics found by BERTopic, as well as the sample overlap with the cluster hierarchy in HDBSCAN of BERTopic and the components found by SCA (see Appendix B.2 for details). Finally, we evaluate the SCA’s transformation into the component space (see Appendix A.2 for details) on the MTEB benchmark (Muennighoff et al., 2023), which consists of various downstream tasks, including sequence classification, summarization and retrieval. We do this to quantify the faithfulness of the representation provided by SCA-components and compare against the base embedding model and a PCA.
@@ 5 Results
Besides the qualitative results in Figure 2, we provide statistical results in Table 1. For all datasets, the method finds a large number of components, with the Trump dataset yielding 182 components in total after merging, the Hausa dataset 331 and the Chinese News dataset 594, surpassing the 55, 102, and 212 respective topics found by BERTopic. These components relate to patterns distinct from the topics found by BERTopic, which is evident from the low average maximum token overlap scores of 1.85, 1.32 and 1.39 on the three datasets respectively. On the Trump dataset, only 8 components with a token overlap of higher than 0.5 are merged, on the Hausa dataset it is 9, and for the Chinese News dataset 19. The average maximum sample overlaps with any of BERTopic’s hierarchical subtopics are small and between 0.077 and 0.093 on all datasets. This indicates that the additional components are not sub- or supertopics of the topics found by BERTopic. The noise rate is close to zero for all datasets, i.e. almost all samples are assigned at least one component. The topic coherence scores are high for all datasets, with the Chinese dataset showing the highest NPMI coherence of 0.120 and the Trump dataset showing the lowest of -0.168, all matching the baseline. The topic diversity is high for all datasets, between 0.7 and 0.8, but about 10-15% lower than in the baseline.
In sum, this shows that the quality of the components found matches that of BERTopic while providing a more nuanced view of the data through the large number of extra components found and the low noise rate. Comparing the results over the hyperparameters on the Trump dataset, we find that for $\alpha>0.7$ and $\mu<0.4$, the method reliably finds a large number of semantic components at least double the number of topics found by BERTopic, usually ranging from three to four times.
@@
The additional components found relate to patterns that intuitively match the expected distribution of the datasets (full Top-10 component tables in Appendix C). For example, in the Trump dataset, beyond the topics about Trump himself, entrepreneurship, and election campaigning found by BERTopic, we uncover components about "fake news", inspiration and role models, paying condolences, and the "witch hunt" narrative among others (see Appendix C.2) underlining SCA’s ability to find important patterns not captured by the baseline. This view is reinforced by the component distributions of single samples, as shown in Figure 1. Similarly, in the Chinese dataset, the additional components relate to important motives in Chinese news reporting, such as protests, family, the trade war with the US, and currency (see Appendix C.3). In the Hausa dataset, the additional insights include components about poverty, courts and law, and work (see Appendix C.4). However, we note, that in this low-resource scenario, some components relate to semantically more narrow concepts, which is possibly due to a lack of expressive power in the base embeddings.
@@
Finally, the results of the topic distribution provided by SCA can match the performance of PCA and the base embedding model in the MTEB benchmark (see Appendix C.5), indicating the method’s faithfulness in representing the information of the samples in the component space.
@@ 6 Conclusion
We introduce Semantic Component Analysis as a novel topic modeling method that can discover multiple semantic components per document. We show that the components found by SCA indeed provide new, more nuanced insights into the data while maintaining a noise rate close to zero and matching the state-of-the-art in terms of topic quality. Furthermore, we show that the method is able to generalize to multiple languages, including a low-resource African language, and to provide a faithful transformation of the embeddings into the component space. We believe that SCA provides a substantial improvement in large-scale text analysis and opens up new possibilities for the interpretation of semantic patterns in short texts.
Limitations
@@
Ethics Statement
@@
Firstly, there is a risk of misinterpretation or misrepresentation of the discovered semantic components. The interpretation of these components relies on the assumption that they represent meaningful and coherent topics or themes in the data. However, there is always a possibility of misinterpreting or misrepresenting the true meaning of these components, leading to incorrect conclusions or biased interpretations. Even earlier in the pipeline, there is a risk of algorithmic bias in the clustering and representation process. Especially the base embedding model can introduce biases based on its training data and model architecture that SCA can inherit. It is important to carefully select and validate the base embedding model to ensure as fair and unbiased representations as possible.
@@
Thirdly, the nature of online social media data, such as Tweets, can introduce additional risks and challenges. Social media data can be subject to manipulation or misinformation, and contain offensive or harmful content. SCA may inadvertently amplify or propagate such content if not properly filtered or evaluated. Additionally, using the insights from SCA might inadvertently lead to amplifying or reinforcing existing biases or stereotypes. It is important to consider the potential impact of the data and the extracted semantic components on individuals, communities, and society as a whole.
@@
Appendices
Appendix A Methodology Appendix
A.1 Hyperparameter Discussion